OpenStudy (anonymous):

Linear Equations. Write the equation in slope-intercept form for a line that passes through the points (-6, 0) and (0, -2).

OpenStudy (anonymous):

slope= (0-(-2))/(-6-0)=2/(-6)=-1/3 y=(-1/3) *(x-(-6)) y=(-1/3)x -2